OpenStudy (anonymous):

Find the inverse of the function: f(x)= cuberoot(x/7)-9

OpenStudy (anonymous):

the inverse means to solve for x .. so lets solve for x add 9 to both sides and you get... y+9=cuberoot of X/7... Now lets get rid of the cube root by multiplying everything to the third root... so (y+9)^3=x/7... Not to get rid of the 7 attached to the x we will multiply by 7.... So for x we get... 7(Y+9)^3=x...... NOw to get our final answer we switch values of x and y simply.... So now the Final answer is Y=(x+9)^3 Hope this helps.

OpenStudy (anonymous):

Ooh I see! But it's still 7(x+9)^3 right because you're still multiplying it by 7?

OpenStudy (anonymous):

Oh yeah oops I forget to include the 7. Yes you are right

OpenStudy (anonymous):

Thanks so much! :)

OpenStudy (anonymous):

You're very welcome!